Required information

Greenwood Company manufactures two products—14,000 units of Product Y and 6,000 units of Product Z. The company uses a plantwide overhead rate based on direct labor-hours. It is considering implementing an activity-based costing (ABC) system that allocates all of its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products Y and Z:

Activity Cost Pool Activity Measure Estimated Overhead Cost Expected
Activity
Machining Machine-hours $ 198,000 10,000 MHs
Machine setups Number of setups $ 86,400 180 setups
Production design Number of products $ 82,000 2 products
General factory Direct labor-hours $ 248,000 10,000 DLHs
Activity Measure Product Y Product Z
Machining 7,800 2,200
Number of setups 50 130
Number of products 1 1
Direct labor-hours 7,800 2,200

5. What is the activity rate for the Product Design activity cost pool?

6. What is the activity rate for the General Factory activity cost pool? (Round your answer to 2 decimal places.)

9. Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Y? (Round your intermediate calculations to 2 decimal places. Round your answers to the nearest dollar amount.)

10. Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Z? (Round your intermediate calculations to 2 decimal places. Round your answers to the nearest dollar amount.)
Answer #1

5) Product design activity rate = 82000/2 = 41000 per product

6) General factory activity = 248000/10000 = 24.80 per hour

9) Total manufacturing overhead assigned to product Y

Product Y
Machining 154440
Machine setup 24000
Product design 41000
General factory 193440
Total overhead 412880

10) Total manufacturing overhead assigned to product Z

Product Z
Machining 43560
Machine setup 62400
Product design 41000
General factory 54560
Total overhead 201520